# Semileptonic decays $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ with the SU(3) flavor analysis

### v1

@@ Abstract
Semileptonic three-body $D\to A\ell^{+}\nu_{\ell}$ decays, nonleptonic $A\to VP$ decays and semileptonic four-body $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ strong or electromagnetic decays are analyzed by the SU(3) flavor symmetry/breaking approach, where $A$ denotes the axial-vector meson, $V$ denotes the vector meson, and $P$ denotes pseudoscalar meson. In terms of the SU(3) flavor symmetry/breaking, the decay amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays or the vertex coefficients of the $A\to VP$ decays are related. The relevant nonperturbative parameters are constrained by the present experimental data. Then the branching ratios of the $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays are predicted by the constrained parameters. We find that some branching ratios are large and they might be measured in the near future experiments. Moreover, the sensitivities of the branching ratios of $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the parameters are also shown, and some decay branching ratios are sensitive to the nonperturbative parameters.
@@ I INTRODUCTION
Semileptonic decays of the charmed mesons are very useful for determining the quark mixing parameters, the values of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements, and searching for new physics beyond the standard model. Among the semileptonic four-body decays $D\to VP\ell^{+}\nu_{\ell}$, only the branching ratios of $D^{+}\to b^{0}_{1}(b_{1}^{0}\to\omega\pi^{0})e^{+}\nu_{e}$ and $D^{0}\to b^{-}_{1}(b_{1}^{-}\to\omega\pi^{-})e^{+}\nu_{e}$ decays have the experiential upper limits PDG2022 . Nevertheless, two semileptonic three-body decays $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ have been found by the CLEO Collaboration CLEO:2007oer and they have been much improved by the BESIII Collaboration BESIII:2019eao ; BESIII:2021uqr . The six nonleptonic decays $K_{1}\to VP$ have been measured by the ACCMOR Collaboration ACCMOR:1981yww . One can use them to study $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays.
@@
For the semileptonic decays, since the leptons do not participate in the strong interaction, the weak and strong dynamics can be separated, the theoretical description of the semileptonic decays are relatively simple. The hadronic transition form factors contain all the strong dynamics in the initial and final hadrons, and they are crucial for testing the theoretical calculations of the involved strong interaction. The structure of the $D\to VP$ form factors is more complicated than for the $D\to V$ or $D\to P$ form factors, nevertheless, we have not found their systematical calculations up to now. In the absence of reliable calculations, the symmetry analysis can provide very useful information about the decays. SU(3) flavor symmetry is independent of the detailed dynamics, it can provide us an opportunity to relate different decay modes. Even though SU(3) flavor symmetry is only an approximate symmetry, it still gives some valuable information about the decays. SU(3) flavor symmetry has been widely used to study hadron decays, including b-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn , c-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Cheng:2012xb ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Muller:2015lua .
@@
The semileptonic $D\to A\ell^{+}\nu_{\ell}$ decays have been studied, for examples, by the QCD sum rules approach Momeni:2020zrb ; Momeni:2019uag ; Zuo:2016msr ; Hu:2021lkl ; Khosravi:2008jw , the light-cone sum rules Huang:2021owr , the quark models Isgur:1988gb ; Scora:1995ty ; Cheng:2017pcq . The nonleptonic $A\to VP$ decays have been studied by many methods, for examples, the chiral unitary approach Geng:2006yb ; Cabrera:2009ep , quark models Ackleh:1996yt ; Bloch:1999vka ; Du:2021zdg , the effective Lagrangian of the strong interaction Lichard:2010zza ; Li:1995aw ; Li:1995tv ; Wess:1967jq ; Roca:2003uk ; Vojik:2010ua . The four-body semileptonic $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays have also been studies by the unitary extensions of chiral perturbation theory Wang:2020pyy . In this work, we will analyze the $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ four-body semileptonic decays by using the SU(3) flavor symmetry/breaking approach. We will first obtain the decay amplitude relations of the $D\to A\ell^{+}\nu_{\ell}$ or the vertex coefficient relations of $A\to VP$ by using the SU(3) flavor symmetry/breaking approach. Then, after ignoring the SU(3) flavor breaking effects due to poor experimental data of $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$, the relevant nonperturbative coefficients and the mixing angles will be constrained by the present experiential data. Finally, the branching ratios, which have not been measured yet, will be predicted by the SU(3) flavor symmetry. In addition, the sensitivities of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$ and $\mathcal{B}(A\to VP)$ decays to the nonperturbative parameters are also displayed.
This paper will be organized as follows. In Sec. II, the results of the semileptonic $D\to A\ell^{+}\nu_{\ell}$ decays will be obtained. In Sec. III, the results of the nonleptonic $A\to VP$ decays will be calculated. The branching ratios of semileptonic four-body $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays will be discussed in Sec. IV. We summarize our results in Sec. V. Finally, all figures and tables are listed in Appendix.
II Semileptonic $D\rightarrow A\ell^{+}\nu_{\ell}$ decays
II.1 Meson multiplets

### v2

# Study of the axial-vector and tensor resonant contributions to the $D\to VP\ell^{+}\nu_{\ell}$ decays based on SU(3) flavor analysis

@@ Abstract
Semileptonic three-body $D\to M\ell^{+}\nu_{\ell}$ decays, non-leptonic $M\to VP$ decays, and semileptonic four-body $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ decays are analyzed using the SU(3) flavor symmetry/breaking approach, where $\ell=e/\mu$, $M=A/T$, and $A/T/V/P$ denote the axial-vector/tensor/vector/pseudoscalar mesons, respectively. In terms of SU(3) flavor symmetry/breaking, the decay amplitudes of the $D\to M\ell^{+}\nu_{\ell}$ decays and the vertex coefficients of the $M\to VP$ decays are related. The relevant non-perturbative parameters of the $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $T\to VP$ decays are constrained by the present experimental data, and the non-perturbative parameters of $D\to T\ell^{+}\nu_{\ell}$ decays are taken from the results in the light-front quark model since no experimental data are available at present. The branching ratios of the $D\to M\ell^{+}\nu_{\ell}$, $M\to VP$, and $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ decays are then predicted. We find that some processes receive both tensor and axial-vector resonant contributions, while other processes receive only axial-vector resonant contributions. In cases where both kinds of resonant contributions exist, the axial-vector contributions are dominant. Some branching ratios with axial-vector resonance states are large, and they may be measured experimentally in the near future. In addition, the sensitivities of the branching ratios of $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the parameters are also investigated, and some decay branching ratios are found to be sensitive to the non-perturbative parameters.
@@ I INTRODUCTION
Semileptonic decays of the charmed mesons are very useful for determining the quark mixing parameters and values of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements, and in searching for new physics beyond the standard model. The intermediate resonances in these semileptonic decays provide a clean environment in which to explore meson spectroscopy, as there is no interference from other particles Ke:2023qzc . Among the semileptonic four-body decays $D\to VP\ell^{+}\nu_{\ell}$, only the branching ratios of $D^{+}\to b^{0}_{1}(b_{1}^{0}\to\omega\pi^{0})e^{+}\nu_{e}$ and $D^{0}\to b^{-}_{1}(b_{1}^{-}\to\omega\pi^{-})e^{+}\nu_{e}$ decays have experimental upper limits PDG2022 . Nevertheless, two semileptonic three-body decays $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ have been found by the CLEO Collaboration CLEO:2007oer , and they have been much improved by the BESIII Collaboration BESIII:2019eao ; BESIII:2021uqr . The six non-leptonic decays $K_{1}\to VP$ were measured by the ACCMOR Collaboration ACCMOR:1981yww . One can use them to study the $D\to VP\ell^{+}\nu_{\ell}$ semileptonic decays with the axial-vector resonance states. As for the $D\to VP\ell^{+}\nu_{\ell}$ decays with the tensor resonance states, only four $T\to VP$ modes have been measured PDG2022 ; as there are no experimental data in either $D\to T\ell^{+}\nu_{\ell}$ or $D\to T(T\to VP)\ell^{+}\nu_{\ell}$ decays, they can be studied by combining theoretical calculations.
@@
For semileptonic decays, since the leptons do not participate in the strong interaction, the weak and strong dynamics can be separated, and the theoretical description of the semileptonic decays is relatively simple. The hadronic transition form factors contain all the strong dynamics in the initial and final hadrons, and they are crucial for testing the theoretical calculations of the involved strong interaction. The structure of the $D\to VP$ form factors is more complicated than for the $D\to V$ or $D\to P$ form factors; nevertheless, we have not found their systematic calculations up to now. In the absence of reliable calculations, symmetry analysis can provide very useful information about the decays. Because SU(3) flavor symmetry is independent of the detailed dynamics, it can provide us an opportunity to relate different decay modes. Even though SU(3) flavor symmetry is only an approximate symmetry, it still provides some valuable information about the decays. SU(3) flavor symmetry has been widely used to study hadron decays, including b-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn and c-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Cheng:2012xb ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Muller:2015lua .
@@
Semileptonic $D\to M\ell^{+}\nu_{\ell}$ decays have been studied, for example, by the quantum chromodynamics (QCD) sum rules approach Momeni:2020zrb ; Momeni:2019uag ; Zuo:2016msr ; Hu:2021lkl ; Khosravi:2008jw , light-cone sum rules Huang:2021owr , and quark models Isgur:1988gb ; Scora:1995ty ; Cheng:2017pcq . The non-leptonic $M\to VP$ decays have been studied by many methods, for examples, the chiral unitary approach Geng:2006yb ; Cabrera:2009ep , quark models Ackleh:1996yt ; Bloch:1999vka ; Du:2021zdg , and the effective Lagrangian of the strong interaction Lichard:2010zza ; Li:1995aw ; Li:1995tv ; Wess:1967jq ; Roca:2003uk ; Vojik:2010ua . The four-body semileptonic $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays have also been studied by the unitary extensions of chiral perturbation theory Wang:2020pyy . In this work, we will analyze the $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ four-body semileptonic decays using SU(3) flavor symmetry/breaking approach. We will first obtain the decay amplitude relations of the $D\to M\ell^{+}\nu_{\ell}$ or the vertex coefficient relations of $M\to VP$ by using SU(3) flavor symmetry/breaking. Then, after ignoring the SU(3) flavor breaking effects due to poor experimental data of $D\to M\ell^{+}\nu_{\ell}$, $M\to VP$ and $D\to M(M\to VP)\ell^{+}\nu_{\ell}$, the relevant non-perturbative coefficients and mixing angles will be constrained by the present experimental data. Finally, the branching ratios which have not yet been measured will be predicted by the SU(3) flavor symmetry. In addition, the sensitivity of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$ and $\mathcal{B}(A\to VP)$ decays to the non-perturbative parameters will be displayed.
This paper is organized as follows. In Sect. II, the $D\rightarrow A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\rightarrow A(A\to PV)\ell^{+}\nu_{\ell}$ decays are presented. In Sect. III, the $D\rightarrow T\ell^{+}\nu_{\ell}$, $T\to VP$ and $D\rightarrow T(T\to PV)\ell^{+}\nu_{\ell}$ decays are discussed. We summarize our results in Sect. IV. Finally, the meson multiplets of the SU(3) flavor group are listed in the Appendix.
II Semileptonic $D\rightarrow A(A\to PV)\ell^{+}\nu_{\ell}$ decays
II.1 Semileptonic $D\rightarrow A\ell^{+}\nu_{\ell}$ decays
II.1.1 Theoretical framework for the $D\to A\ell^{+}\nu_{\ell}$ decays
$\mathcal{M}(D\rightarrow A\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}\sum_{m,n}g_{mn}L^{\lambda_{\ell}\lambda_{\nu}}_{m}H^{\lambda_{A}}_{n}.$ (2)
Leptonic amplitudes $L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ and hadronic amplitudes $H^{\lambda_{A}}_{n}$ are represented as
$H^{\lambda_{A}}_{n}$ $=$ $V_{cq}^{*}\epsilon_{\beta}^{*}(n)\langle{A}(p,\varepsilon^{*})|\bar{q}\gamma^{\beta}(1-\gamma_{5})c|D(p_{D})\rangle,$ (4)
where $\lambda_{A}=0,\pm 1$, $\lambda_{\ell}=\pm\frac{1}{2}$, and $\lambda_{\nu}=+\frac{1}{2}$ are the particle helicity, and $\epsilon(m)$ with $m=0,t,\pm 1$ and $\varepsilon^{*}$ are the polarization vectors of the virtual $W$ and axial-vector mesons $A$, respectively.
$A(p,\varepsilon^{*})\left|\bar{q}\gamma_{\mu}(1-\gamma_{5})c\right|D(p_{D})\right>$ $=$ $\frac{2iA(q^{2})}{m_{D}-m_{A}}\epsilon_{\mu\nu\alpha\beta}\varepsilon^{*\nu}p^{\alpha}_{D}p^{\beta}$ (5)
$+q_{\mu}(\varepsilon^{*}.p_{D})\frac{2m_{A}}{q^{2}}[V_{3}(q^{2})-V_{0}(q^{2})],$
where $q^{2}\equiv(p_{D}-p)^{2}$, $A(q^{2})$ and $V_{0,1,2,3}(q^{2})$ are the form factors of the $D\to A$ transition, and $m_{A,D,\ell}$ are the masses of $A,D,\ell$. Ten hadronic amplitudes can then be written as Sun:2011ssd ; Li:2009tx ; Colangelo:2019axi
$H_{t}$ $=$ $\frac{2m_{D}|\vec{p}_{A}|}{\sqrt{q^{2}}}V_{0}(q^{2}),$ (8)
The differential branching ratios are Ivanov:2019nqd
where $\lambda\equiv\lambda(m_{D}^{2},m_{A}^{2},q^{2})$, $m_{\ell}^{2}\leq q^{2}\leq(m_{D}-m_{A})^{2}$, and $\tau_{D}$ are the mean lifetime of $D$ meson. The branching ratios are usually obtained by Eq. (9) and the form factors; however, the results of the form factors are dependent on different approaches.